Vanderwaals equation is given by [P+a/V2](v-b)=RT
Where P is pressure exerted by a gas.T is absolute temperature,V is volume of a gas and a and b are constants.Find dimensions and units of a and b

Grade:12

2 Answers

Arun
4 years ago
Vander Waal's Universal equation for a non ideal gas :
 
 ( P + a/v² ) * (V - b) = n R T
 
with the symbols having the given meanings.
 
P = pressure. V = volume. R = Universal Gas constant. T = absolute temperature. n = number of moles of the gas.
 
Since P has dimensions of Force/Area, a/V² also will have the same units and dimensions.
 
P = [ M ] [ L⁻¹ ] [ T⁻² ] = a / V² = a [ L³ ]⁻²
 
So constant a = [ M ] [ L⁵ ] [ T⁻² ]
 
Since b has same units and dimensions as that of V,
 
  b = [ L³ ]
 
Hope it helps
Vikas TU
4 years ago
Dear student 
Please refer the below link 
Good Luck 
Cheers